# Quantum-Adversary-Resilient Evidence Structures and Migration Strategies for Regulated AI Audit Trails

### v1

@@ Abstract
Constant-size cryptographic evidence records are increasingly used to build audit trails for regulated AI workloads in clinical, pharmaceutical, and financial settings, where each execution is summarized by a compact, verifiable record of code identity, model version, data digests, and platform measurements. Existing instantiations, however, typically rely on classical signature schemes whose long-term security is threatened by quantum-capable adversaries. In this paper we formalize security notions for evidence structures in the presence of quantum adversaries and study post-quantum (PQ) instantiations and migration strategies for deployed audit logs. We recall an abstraction of constant-size evidence structures and introduce game-based definitions of Q-Audit Integrity, Q-Non-Equivocation, and Q-Binding, capturing the inability of a quantum adversary to forge, equivocate, or rebind evidence items. We then analyze a hash-and-sign instantiation in the quantum random-oracle model (QROM), assuming an existentially unforgeable PQ signature scheme against quantum adversaries, and show that the resulting evidence structure satisfies these notions under standard assumptions. Building on this, we present three migration patterns for existing evidence logs—hybrid signatures, re-signing of legacy evidence, and Merkle-root anchoring—and analyze their security, storage, and computational trade-offs. A case study based on an industrial constant-size evidence platform for regulated AI at Codebat Technologies Inc. suggests that quantum-safe audit trails are achievable with moderate overhead and that systematic migration can significantly extend the evidentiary lifetime of existing deployments.
@@ 1 Introduction
Audit trails play a central role in regulated environments where automated computation affects safety- or compliance-critical decisions. In domains such as clinical AI, pharmaceutical development, and high-stakes financial analytics, regulators require that organizations be able to explain and justify how a particular result was produced: which code and model were used, which data were processed, under which policies, on which platform, and at what time. Recent system designs address these needs by representing each execution as a constant-size cryptographic evidence item that compactly captures the essential provenance and integrity information and can be stored and verified at scale (e.g., [kao2025constant, securelogging, transparency]), and recent work has begun to study quantum-resilient designs for distributed log storage itself [somarapu2022qrlogs].
@@
In this work we address these gaps. We build on an abstraction of constant-size evidence structures introduced in prior work [kao2025constant], which models a system where each event in a workflow is associated with a fixed-length evidence record produced by a Generate algorithm and consumed by a Verify algorithm. Our focus here is on extending this abstraction to a quantum-adversary setting, providing both formal definitions and practically oriented migration strategies.
@@
    Post-quantum instantiations in the QROM. We analyze a hash-and-sign instantiation of constant-size evidence structures in the quantum random-oracle model, assuming an underlying post-quantum signature scheme that is existentially unforgeable under chosen-message attacks (EUF-CMA) against quantum adversaries. We prove that under standard assumptions the instantiated scheme satisfies our quantum-adversary security notions.
@@
    Industrial case study. We outline a case study using a constant-size evidence system implemented in an industrial platform for regulated AI workloads at Codebat Technologies Inc. We discuss how the proposed migration patterns map to realistic deployment constraints and estimate the cost of securing large evidence corpora.
The rest of the paper is organized as follows. Section 2 recalls the abstraction of constant-size evidence structures and reviews quantum-adversary and PQ primitive models. Section 3 introduces our quantum-adversary security notions. Section 4 presents and analyzes a post-quantum instantiation. Section 5 develops migration patterns for deployed evidence logs. Section 6 discusses an industrial case study. Section 7 reviews related work, and Sections 89 conclude with open problems and future directions.
@@
Compared to classical secure logging and transparency systems (e.g., [securelogging, forwardsecurelogging, haberstornetta1991, merkle1989certified, ct, transparency]), our work treats constant-size evidence structures as the primary cryptographic object and explicitly analyzes their security against quantum adversaries in the QROM. Existing PQC work on logging and auditing (e.g., [somarapu2022qrlogs, khan2025pqblockchainaudit, yusuf2025qrlogintegrity]) typically either focuses on blockchain-based storage or on specific application domains. In contrast, we provide (i) generic quantum-adversary security notions for evidence structures, and (ii) migration patterns that are designed to retrofit already-deployed constant-size evidence logs in regulated AI systems.
@@ 2.1 Constant-Size Evidence Structures
We briefly recall the abstraction of constant-size evidence structures introduced in [kao2025constant], which we treat here as a black-box building block.
@@
The original work [kao2025constant] focuses on classical adversaries, defining integrity and linkage properties and providing a concrete instantiation where each $\mathsf{Ev}$ is a 288- or 320-byte record. In this paper we treat $\mathsf{Generate}$ and $\mathsf{Verify}$ as abstract interfaces that we instantiate with post-quantum primitives and analyze against quantum adversaries.
@@ 2.2 Quantum Adversaries and Post-Quantum Primitives
We adopt the quantum random-oracle model (QROM) for hash functions $H:\{0,1\}^{*}\rightarrow\{0,1\}^{\lambda}$, following the treatment of Zhandry and subsequent work [zhandry2012qrom, bonehzhandry2013qro]. In the QROM, $H$ is modeled as a uniformly random function, and the adversary may query it in quantum superposition. We assume that $H$ is collision-resistant against quantum adversaries, i.e., no QPT adversary can find distinct $x,x^{\prime}$ with $H(x)=H(x^{\prime})$ with non-negligible probability.
@@
We consider a signature scheme $\Sigma=(\mathsf{KeyGen}^{\Sigma},\mathsf{Sign}^{\Sigma},\mathsf{Verify}^{\Sigma})$ with standard syntax:
@@
Concretely, $\Sigma$ may correspond to a lattice-based or hash-based scheme such as those standardized or under consideration in post-quantum cryptography efforts (e.g., [nistpqc, unruh2017pqsignatures]), but in this paper we treat it as an abstract primitive satisfying the above property. Real-world deployments face non-trivial challenges around key sizes, performance, and migration, as surveyed in [tan2022pqsignsurvey].
@@ 3 Security Notions Against Quantum Adversaries
3.1 System Model with Quantum Adversaries
We consider an environment where an evidence generator uses $\mathsf{Generate}$ to produce evidence items for events and stores them in an append-only audit log. An auditor uses $\mathsf{Verify}$ and (optionally) $\mathsf{Link}$ to check evidence for consistency with a claimed workflow.
@@ 3.2 Game-Based Definitions
    • for both $b\in\{0,1\}$, all pairs in $\mathcal{L}^{(b)}$ verify under $\mathsf{Verify}(\mathsf{pp},\mathsf{pk},\cdot,\cdot)$; and
    • there exists at least one index $i$ such that $(e^{(0)}_{i},\mathsf{Ev}^{(0)}_{i})\neq(e^{(1)}_{i},\mathsf{Ev}^{(1)}_{i})$ but both are presented at the same log position $i$ in their respective views.
@@
    The experiment outputs $1$ if both bullets hold.
@@
The second bullet captures the notion that the two logs represent conflicting stories about the same workflow (e.g., a particular event is associated with different evidence items), while both appear individually valid.
@@ 3.3 Relation to Classical Notions
In the special case where adversaries are classical and have only classical access to the hash oracle, our definitions reduce to classical notions of integrity and non-equivocation for secure logging and transparency systems [securelogging, ct]. Our experiments focus on the inability to create new valid evidence, to maintain two conflicting yet valid views of the log, and to rebind an evidence item. These correspond to the informal goals discussed in earlier work on constant-size evidence structures in a classical setting (e.g., [kao2025constant]) but here explicitly account for quantum-accessible oracles and quantum EUF-CMA adversaries.
@@ 4.1 Hash-and-Sign Construction in the QROM
We assume that for each event $e$ we can deterministically compute a canonical encoding $\mathsf{Fields}(e)\in\{0,1\}^{*}$ that concatenates all fields to be included in the evidence record (e.g., code hash, input digest, platform measurement, timestamps). The precise layout of these fields follows the constant-size design from [kao2025constant]; we abstract it here as a bitstring.
@@ 4.2 Post-Quantum Signature Choices
From the perspective of constant-size evidence records, the signature size directly contributes to the overall record size. If the base layout for $x$ is, say, $160$ bytes (including digests and metadata) and the signature is $128$ bytes, the resulting record size is $288$ bytes (plus any padding or alignment required), which is in the same range as prior designs [kao2025constant]. The throughput impact can be estimated by benchmarking the chosen $\Sigma$ on representative hardware; our case study in Section 6 sketches such an analysis.
@@ 4.3 Security Theorems
We now state our main security theorem for the hash-and-sign instantiation. Proofs follow standard reduction techniques from signature and QROM analyses; we provide sketches here and defer full details to an extended version.
Theorem 1 (Q-Audit Integrity and Q-Binding).
Assume that $H$ is collision-resistant against quantum adversaries in the QROM and that $\Sigma$ is EUF-CMA secure against quantum adversaries. Then the hash-and-sign evidence structure described above satisfies Q-Audit Integrity and Q-Binding in the sense of the experiments $\mathsf{Exp}^{\mathsf{qaudit}}$ and $\mathsf{Exp}^{\mathsf{qbind}}$.
Proof sketch.
We outline the argument for Q-Audit Integrity; the Q-Binding property is similar.
Suppose there exists a QPT adversary $\mathcal{A}$ that wins $\mathsf{Exp}^{\mathsf{qaudit}}_{\mathcal{A}}(1^{\lambda})$ with non-negligible probability $\epsilon(\lambda)$. We construct a QPT adversary $\mathcal{B}$ that breaks either the EUF-CMA security of $\Sigma$ or the collision resistance of $H$.
Adversary $\mathcal{B}$ is given a public key $\mathsf{pk}^{\Sigma}$ and oracle access to $\mathsf{Sign}^{\Sigma}$ and $H$ in the QROM. It sets $\mathsf{pk}=\mathsf{pk}^{\Sigma}$ and simulates the $\mathsf{Generate}$ oracle for $\mathcal{A}$ by: on input $e$, computing $x=\mathsf{Fields}(e)$, $m=H(x)$, and obtaining $\sigma\leftarrow\mathsf{Sign}^{\Sigma}(m)$ from its signing oracle, then returning $\mathsf{Ev}=(x,\sigma)$. It maintains the set $\mathcal{Q}$ of generated pairs as in the experiment. Note that $\mathcal{B}$ can answer all of $\mathcal{A}$’s hash queries by forwarding them to the QROM oracle.
Eventually, $\mathcal{A}$ outputs $(e^{\star},\mathsf{Ev}^{\star})$ with $\mathsf{Ev}^{\star}=(x^{\star},\sigma^{\star})$ such that $\mathsf{Verify}(\mathsf{pp},\mathsf{pk},e^{\star},\mathsf{Ev}^{\star})=1$ and $(e^{\star},\mathsf{Ev}^{\star})\notin\mathcal{Q}$. There are two cases:
@@
    If $x^{\star}$ equals $\mathsf{Fields}(e)$ for some event $e$ for which $\mathcal{B}$ already obtained a signature via the signing oracle, then $(e^{\star},\mathsf{Ev}^{\star})\in\mathcal{Q}$, contradicting the success condition of $\mathcal{A}$. Thus, this case can only occur with negligible probability due to simulation inconsistencies, which can be bounded using standard QROM techniques.

### v2

# Post-Quantum-Resilient Audit Evidence for Long-Lived Regulated Systems: Security Models, Migration Patterns, and Case Study

@@ Abstract
Constant-size cryptographic evidence records are increasingly used to build audit trails for regulated AI workloads in clinical, pharmaceutical, and financial settings, where each execution is summarized by a compact, verifiable record of code identity, model version, data digests, and platform measurements. Existing instantiations, however, typically rely on classical signature schemes whose long-term security is threatened by quantum-capable adversaries. In this paper we formalize security notions for evidence structures in the presence of quantum adversaries and study post-quantum (PQ) instantiations and migration strategies for deployed audit logs. We recall an abstraction of constant-size evidence structures and introduce game-based definitions of Q-Audit Integrity, Q-Non-Equivocation, and Q-Binding, capturing the inability of a quantum adversary to forge, equivocate, or rebind evidence items. We then analyze a hash-and-sign instantiation in the quantum random-oracle model (QROM), assuming an existentially unforgeable PQ signature scheme against quantum adversaries, and prove via tight reductions that the resulting evidence structure satisfies all three notions under standard assumptions (collision-resistant hashing and quantum EUF-CMA signatures). Building on this, we present three migration patterns for existing evidence logs—hybrid signatures, re-signing of legacy evidence, and Merkle-root anchoring—and analyze their security, storage, and computational trade-offs. A case study based on an industrial constant-size evidence platform for regulated AI at Codebat Technologies Inc., implemented in Rust and benchmarked on commodity hardware, suggests that quantum-safe audit trails are achievable with moderate overhead and that systematic migration can significantly extend the evidentiary lifetime of existing deployments.
@@ 1 Introduction
Audit trails play a central role in regulated environments where automated computation affects safety- or compliance-critical decisions. In domains such as clinical AI, pharmaceutical development, and high-stakes financial analytics, regulators require that organizations be able to explain and justify how a particular result was produced: which code and model were used, which data were processed, under which policies, on which platform, and at what time. Recent system designs address these needs by representing each execution as a constant-size cryptographic evidence item that compactly captures the essential provenance and integrity information and can be stored and verified at scale (e.g., [7, 14, 9]), and recent work has begun to study quantum-resilient designs for distributed log storage itself [16].
@@
In this work we address these gaps. We build on an abstraction of constant-size evidence structures introduced in prior work [7], which models a system where each event in a workflow is associated with a fixed-length evidence record produced by a Generate algorithm and consumed by a Verify algorithm. Our focus here is on extending this abstraction to a quantum-adversary setting, providing both formal definitions and practically oriented migration strategies.
@@
    Post-quantum instantiations with complete proofs in the QROM. We analyze a hash-and-sign instantiation of constant-size evidence structures in the quantum random-oracle model and provide three complete reduction proofs with explicit advantage bounds, showing that Q-Audit Integrity, Q-Binding, and Q-Non-Equivocation all hold under collision-resistant hashing and quantum EUF-CMA signatures.
@@
    Industrial case study with PQ benchmarks. We outline a case study using a constant-size evidence system implemented in an industrial platform for regulated AI workloads at Codebat Technologies Inc., including concrete performance comparisons across post-quantum signature schemes (ML-DSA, SLH-DSA) to quantify migration overhead.
  5. 5.
    Systematic comparison. We provide a feature-level comparison with existing post-quantum logging and audit systems, identifying the gaps that our work addresses.
The rest of the paper is organized as follows. Section 2 recalls the abstraction of constant-size evidence structures and reviews quantum-adversary and PQ primitive models. Section 3 describes the deployment context and constant-size evidence layout in our industrial platform. Section 4 introduces our quantum-adversary security notions. Section 5 presents and analyzes a post-quantum instantiation. Section 6 develops migration patterns for deployed evidence logs. Section 7 discusses an industrial case study. Section 8 reviews related work, and Sections 910 conclude with open problems and future directions.
@@
Compared to classical secure logging and transparency systems [14, 3, 6, 11, 10, 9], our work treats constant-size evidence structures as the primary cryptographic object and explicitly analyzes their security against quantum adversaries in the QROM. Existing PQC work on logging and auditing (e.g., [16, 8, 19]) typically either focuses on blockchain-based storage or on specific application domains. In contrast, we provide (i) generic quantum-adversary security notions for evidence structures, and (ii) migration patterns that are designed to retrofit already-deployed constant-size evidence logs in regulated AI systems.
@@ 2.1 Constant-Size Evidence Structures
We briefly recall the abstraction of constant-size evidence structures introduced in [7], which we treat here as a black-box building block.
@@
The original work [7] focuses on classical adversaries, defining integrity and linkage properties and providing a concrete instantiation where each $\mathsf{Ev}$ is a 288- or 320-byte record. In this paper we treat $\mathsf{Generate}$ and $\mathsf{Verify}$ as abstract interfaces that we instantiate with post-quantum primitives and analyze against quantum adversaries.
@@ 2.2 Quantum Adversaries and Post-Quantum Primitives
We adopt the quantum random-oracle model (QROM) for hash functions $H:\{0,1\}^{*}\rightarrow\{0,1\}^{\lambda}$, following the treatment of Zhandry and subsequent work [20, 5]. In the QROM, $H$ is modeled as a uniformly random function, and the adversary may query it in quantum superposition. We assume that $H$ is collision-resistant against quantum adversaries, i.e., no QPT adversary can find distinct $x,x^{\prime}$ with $H(x)=H(x^{\prime})$ with non-negligible probability.
@@
We consider a signature scheme $\Sigma$ with standard syntax:
@@
Concretely, $\Sigma$ may correspond to a lattice-based or hash-based scheme such as those standardized or under consideration in post-quantum cryptography efforts (e.g., [12, 18]), but in this paper we treat it as an abstract primitive satisfying the above property. Real-world deployments face non-trivial challenges around key sizes, performance, and migration, as surveyed in [17].
@@
3 System Context: Constant-Size Evidence for Regulated AI
Before introducing our quantum-adversary security model, we describe the deployment context and the constant-size evidence structure that motivates our formal treatment. This section draws on our prior work on constant-size evidence packs for regulated AI [7] and summarizes the key system design choices relevant to quantum-resilient migration.
3.1 Deployment Scenario
We consider regulated AI deployments in domains such as clinical decision support, pharmaceutical development, and financial analytics, where each AI workload execution must be auditable for compliance and forensic purposes. In a typical deployment, the platform orchestrates AI jobs across a cluster of container-based trusted execution environments (TEEs). For each job, the runtime collects attestation and provenance attributes and produces a constant-size evidence record that is appended to an audit log.
Unlike short-lived session keys in TLS, audit evidence for regulated AI workloads is meant to support long-term accountability and forensic investigations. Hospitals, financial institutions, and pharmaceutical companies routinely retain audit logs for ten years or longer due to regulatory requirements. This creates an attractive “harvest-now, forge-later” target: an adversary can exfiltrate classical evidence logs today and exploit future quantum capabilities to retroactively forge or repudiate past AI executions.
3.2 Evidence Record Layout
Each AI job produces a constant-size evidence record capturing the key attestation and provenance attributes needed for regulated audit trails. Table 1 summarizes a representative layout used in our industrial deployment at Codebat Technologies Inc.
Table 1: Representative constant-size evidence layout in our regulated AI deployment, following the abstraction from [7]. Each field is a $\lambda$-bit hash output (typically $\lambda=256$).
Field Description Size (bits)
$f_{0}$: context Workflow/job identifier commitment 256
$f_{1}$: model Hash of model binary / parameters 256
$f_{2}$: code Hash of container image / code snapshot 256
$f_{3}$: input Hash of input dataset or configuration 256
$f_{4}$: output Hash of output artifacts 256
$f_{5}$: policy Policy state / IRB identifier commitment 256
$f_{6}$: environment TEE measurement / platform attestation 256
$f_{7}$: link Link to previous evidence or audit chain 256
$\sigma$: signature Ed25519 or PQ signature 512–2048
Total ($k=8$ fields + signature) 320–512 bytes
The fixed-size layout ensures that each evidence record occupies the same storage footprint regardless of the complexity of the underlying AI job, enabling predictable storage costs and $O(1)$ per-record verification. The signature field currently uses a classical scheme (e.g., ECDSA or Ed25519) but is designed to accommodate post-quantum signatures as part of the migration strategies discussed in Section 6.
3.3 Implementation and Baseline Performance
We have implemented the constant-size evidence generator in Rust as part of an industrial platform for regulated AI deployments [7]. The implementation uses a standard 256-bit hash function (SHA-256) and an Edwards-curve-based signature scheme (Ed25519). The number of fields $k$ is fixed at system setup time; in practice, $k\in[8,12]$ is sufficient for most clinical and pharmaceutical workflows.
[…]
Table 2: Evidence generation and verification throughput on commodity hardware.
[…]
Generation (single-threaded) $3.5\times 10^{4}$ 28.4 $k$ fixed, $\lambda=256$
Generation (16 CPU cores) $2.8\times 10^{5}$ 5.7 near-linear scaling
Verification (16 CPU cores) $2.5\times 10^{5}$ 6.1 batch of $10^{5}$ items
Verification (GPU) $4.0\times 10^{5}$ 2.5 batch of $10^{6}$ items
The fixed-size layout of evidence items leads to regular memory access patterns and uniform control flow, which are favorable for both vectorized CPU execution and GPU kernels. This $O(1)$ per-event cost is crucial for long-term retention, as regulated environments are expected to keep audit logs for 10–30 years while enforcing strict storage and verification latency budgets. A deployment handling $10^{3}$ AI jobs per day accumulates on the order of $10^{6}$$10^{7}$ evidence records over a decade, all of which must remain efficiently verifiable throughout their retention period.
@@ 4 Security Notions Against Quantum Adversaries
4.1 System and Threat Model
We consider a regulated AI deployment where each execution of an AI workload produces a constant-size evidence record, as described in Section 3, and these records are appended to an audit log that is expected to be stored and queried over many years. The evidence generator is integrated with the AI platform’s scheduling and container orchestration layer, while one or more auditors rely on the evidence log to reconstruct and verify past executions during compliance checks and incident response.
Formally, an evidence generator uses $\mathsf{Generate}$ to produce evidence items for events and stores them in an append-only audit log. An auditor uses $\mathsf{Verify}$ and (optionally) $\mathsf{Link}$ to check evidence for consistency with a claimed workflow.
@@ 4.2 Game-Based Definitions
    • for both $b\in\{0,1\}$, all pairs in $\mathcal{L}^{(b)}$ verify under $\mathsf{Verify}(\mathsf{pp},\mathsf{pk},\cdot,\cdot)$;
    • if $\mathsf{Link}$ is defined, $\mathsf{Link}(\mathsf{pp},\mathcal{L}^{(0)})=\mathsf{Link}(\mathsf{pp},\mathcal{L}^{(1)})$; and
    • there exists at least one index $i$ such that $(e^{(0)}_{i},\mathsf{Ev}^{(0)}_{i})\neq(e^{(1)}_{i},\mathsf{Ev}^{(1)}_{i})$.
@@
    The experiment outputs $1$ if all conditions hold.
@@
The second and third bullets together capture the notion that the adversary presents two conflicting but individually valid views of the same workflow: the logs agree on their compact summary (linkage proof) yet disagree about what actually happened at some step. Without the linkage condition, the adversary could trivially win by reordering honestly generated items; the linkage mechanism (e.g., hash chains) prevents such reordering.
@@ 4.3 Relation to Classical Notions
In the special case where adversaries are classical and have only classical access to the hash oracle, our definitions reduce to classical notions of integrity and non-equivocation for secure logging and transparency systems [14, 10]. Our experiments focus on the inability to create new valid evidence, to maintain two conflicting yet valid views of the log, and to rebind an evidence item. These correspond to the informal goals discussed in earlier work on constant-size evidence structures in a classical setting (e.g., [7]) but here explicitly account for quantum-accessible oracles and quantum EUF-CMA adversaries.
@@ 5.1 Hash-and-Sign Construction in the QROM
We assume that for each event $e$ we can deterministically compute a canonical encoding $\mathsf{Fields}(e)$ (a bitstring in $\{0,1\}^{*}$) that concatenates all fields to be included in the evidence record (e.g., code hash, input digest, platform measurement, timestamps). The precise layout of these fields follows the constant-size design from [7]; we abstract it here as a bitstring.
@@ 5.2 Post-Quantum Signature Choices
From the perspective of constant-size evidence records, the signature size directly contributes to the overall record size. If the base layout for $x$ is, say, $160$ bytes (including digests and metadata) and the signature is $128$ bytes, the resulting record size is $288$ bytes (plus any padding or alignment required), which is in the same range as prior designs [7]. The throughput impact can be estimated by benchmarking the chosen $\Sigma$ on representative hardware; our case study in Section 7 sketches such an analysis.
5.3 Hash-Chain Linkage
We instantiate the $\mathsf{Link}$ algorithm using a hash-chain construction. Fix an initialization vector $\ell_{0}=0^{\lambda}$. For a sequence of evidence records $(\mathsf{Ev}_{1},\dots,\mathsf{Ev}_{T})$, define
$\ell_{i}\;=\;H(\ell_{i-1}\;\|\;\mathsf{Ev}_{i}),\quad i=1,\dots,T,$
and set $\mathsf{Link}(\mathsf{pp},\mathsf{Ev}_{1},\dots,\mathsf{Ev}_{T})=\ell_{T}$. The chain tip $\ell_{T}$ serves as a compact summary of the entire log: any modification, reordering, or omission of evidence items changes $\ell_{T}$ with overwhelming probability under the collision resistance of $H$. Auditors compare chain tips to detect inconsistencies between different views of the same log.
@@ 5.4 Security Theorems
We now state and prove the security of the hash-and-sign instantiation with respect to the quantum-adversary notions introduced in Section 4. We provide three theorems corresponding to the three security properties, each with an explicit advantage bound and a complete reduction proof.
Theorem 1 (Q-Audit Integrity).
Let $\mathcal{E}$ be the hash-and-sign evidence structure defined above, instantiated with a PQ signature scheme $\Sigma$ and hash function $H$ in the QROM. For any QPT adversary $\mathcal{A}$ making at most $q_{s}$ queries to the $\mathsf{Generate}$ oracle and at most $q_{H}$ quantum queries to $H$, there exist QPT adversaries $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ such that
$\mathsf{Adv}^{\mathsf{qaudit}}_{\mathcal{E}}(\mathcal{A},\lambda)\;\leq\;\mathsf{Adv}^{\mathrm{euf\text{-}cma}}_{\Sigma}(\mathcal{B}_{1},\lambda)\;+\;\mathsf{Adv}^{\mathrm{cr}}_{H}(\mathcal{B}_{2},\lambda),$
where $\mathcal{B}_{1}$ makes at most $q_{s}$ signing queries and $q_{H}$ hash queries, and $\mathcal{B}_{2}$ makes at most $q_{H}$ hash queries.
Proof.
Suppose $\mathcal{A}$ wins $\mathsf{Exp}^{\mathsf{qaudit}}_{\mathcal{A}}(1^{\lambda})$ with non-negligible advantage $\epsilon(\lambda)$. We construct reductions to the EUF-CMA security of $\Sigma$ and the collision resistance of $H$.
Reduction $\mathcal{B}_{1}$ (to EUF-CMA of $\Sigma$). $\mathcal{B}_{1}$ receives a challenge public key $\mathsf{pk}^{\Sigma}$ and access to a signing oracle $\mathsf{Sign}^{\Sigma}(\mathsf{sk}^{\Sigma},\cdot)$ and the quantum random oracle $H$. It sets $\mathsf{pk}=\mathsf{pk}^{\Sigma}$ and simulates the experiment for $\mathcal{A}$:
  • Hash queries. $\mathcal{B}_{1}$ forwards all of $\mathcal{A}$’s quantum hash queries to the QROM oracle $H$.
  • $\mathsf{Generate}$ queries. On input event $e$, $\mathcal{B}_{1}$ computes $x=\mathsf{Fields}(e)$ and $m=H(x)$, obtains $\sigma\leftarrow\mathsf{Sign}^{\Sigma}(\mathsf{sk}^{\Sigma},m)$ from its signing oracle, returns $\mathsf{Ev}=(x,\sigma)$, and records $(e,\mathsf{Ev})$ in $\mathcal{Q}$ and $m$ in a set $\mathcal{M}$ of signed messages.
When $\mathcal{A}$ outputs $(e^{\star},\mathsf{Ev}^{\star})$ with $\mathsf{Ev}^{\star}=(x^{\star},\sigma^{\star})$ such that $\mathsf{Verify}(\mathsf{pp},\mathsf{pk},e^{\star},\mathsf{Ev}^{\star})=1$ and $(e^{\star},\mathsf{Ev}^{\star})\notin\mathcal{Q}$, $\mathcal{B}_{1}$ computes $m^{\star}=H(x^{\star})$ and considers two cases:
@@
    Case 1: $m^{\star}\notin\mathcal{M}$. Then $(m^{\star},\sigma^{\star})$ is a valid signature on a message never submitted to the signing oracle, constituting a forgery against $\Sigma$. $\mathcal{B}_{1}$ outputs $(m^{\star},\sigma^{\star})$.